OpenStudy (anonymous):

Anyone know how to solve the following equation: -2/3x+3/4=1/3

OpenStudy (anonymous):

1. add 3/4 to both sides of the equation. 2. divide both sides by -2/3

OpenStudy (anonymous):

wait, subtract not add the first step

OpenStudy (anonymous):

-8x + 9 = 4 -8x = -5 x = 5/8

OpenStudy (anonymous):

(2/(3x+3)/4)=1/3 time flip (2*4)/(3x+3)=1/3 (8/(3x+3)=1/3 cross mutiplfy 3x+3=24 3x= 24-3 x = 21/3 x= 7

OpenStudy (anonymous):

okay the way this is written out is different then you have it nancy

OpenStudy (anonymous):

thank

OpenStudy (anonymous):

there are no parenthesis

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but in here I write parenthesi because I don't know how to write another way but t think my answer is corect, see the last part i didn't write parenthesi

OpenStudy (anonymous):

If you follow order of operations, thats not the right way to solve it.

OpenStudy (anonymous):

OK , i THINK i MISS UBDERSTAND ,THANK

OpenStudy (anonymous):

x = 5/8

OpenStudy (anonymous):

It seems that x=5/8 is the answer because when I plug in 5/8, the answer works out. Thanks y'all. :)